Grade 11Mechanics

A force F=AI+bj+ck is acted upon a body of mass m .if the body starts from rest and was at the origin initially . it new co-ordinate after time t is

Devesh
9 Years agoGrade 11

2 Answers

Indrajit dodiya
9 Years ago
Acceleration a=(A/m)i + (b/m)j +(c/m)kNow , second equation of motion, s=ut + 1/2 at2 for all u=0So new co-ordinates (At2/2m , bt2/2m , ct2/2m)
Spring
5 Years ago
F=ma
a=F/m
a=(a/m)I+(b/m)j+(c/m)k
S=ut+1/2at^2
S=0+1/2((a/m)I+(b/m)j+(c/m)k)t^2
S=at^2/2m+bt^2/2m+ct^2/2m